OpenStudy (anonymous):

can u cancel out x^2? problem= [(x^2)+2x]/(x^2)-x

OpenStudy (anonymous):

may be canceled if you were multiplying

OpenStudy (anonymous):

but im dividing woud it still work

OpenStudy (ashwinram):

\[\frac{x^{2}+2x}{x^{2}}-x\]

OpenStudy (anonymous):

not to say that x ^ 2 is not accompanied by x ^ 2 (+) plus sign. example (x ^ 2) 3 / (x ^ 2) 2y = 3/2y

OpenStudy (ashwinram):

If you have: (a + b) / a, you have to distribute the /a over both the numerators. So, it's (a/a) + (b/a). The first a/a cancel out (you get 1) but you still have the second b/a. So it's 1 + (b/a).

OpenStudy (ashwinram):

Likewise, \[(x ^{2} + 2x)/x ^{2}\] will give you \[(x ^{2}/x ^{2}) + (2x/x ^{2})\], and now you simplify.

OpenStudy (anonymous):

\[(x ^{2}+2x)\div(x ^{2}-x)\] is the equation. how would at be simplified?